### v1

# Diffeomorphic approximation of piecewise affine homeomorphisms

Abstract.
Given any $f$ a locally finitely piecewise affine homeomorphism of $\Omega\subset\mathbb{R}^{d}$ onto $\Delta\subset\mathbb{R}^{d}$ (for $d=3,4$) such that $f\in W^{1,p}(\Omega,\mathbb{R}^{d})$ and $f^{-1}\in W^{1,q}(\Delta,\mathbb{R}^{d})$, $1\leq p,q<\infty$ and any $\varepsilon>0$ we construct a diffeomorphism $\tilde{f}$ such that
$\|f-\tilde{f}\|_{W^{1,p}(\Omega,\mathbb{R}^{d})}+\|f^{-1}-\tilde{f}^{-1}\|_{W^{1,q}(\Delta,\mathbb{R}^{d})}<\varepsilon.$
Key words and phrases:
diffeomorphic approximation, Sobolev approximation, piecewise affine homeomorphisms
2010 Mathematics Subject Classification:
Primary 41A29; Secondary 41A30, 57Q55
The first author was supported by the grant GAČR P201/24-10505S and by the Ministry of Education, Youth and Sport of the Czech Republic grant number LL2105 CONTACT. The second author was supported by GNAMPA and by the European Union - NextGenerationEU within the framework of PNRR Mission 4 - “PRIN 2022” - grant number 2022BCFHN2 - Advanced theoretical aspects in PDEs and their applications CUP I53D23002300006, the third author was supported by the Charles University Research Center program No. UNCE/24/SCI/022 and the project SVV-2025-260837-.
1. Introduction
The approximation of homeomorphisms by diffeomorphisms represents a fundamental problem in geometric analysis and differential topology. In particular, understanding when and how homeomorphisms with limited regularity can be approximated by piecewise linear homeomorphisms, smooth homeomorphisms or diffeomorphisms, while preserving key analytic properties has significant implications across several areas of mathematics including geometric function theory, partial differential equations, and continuum mechanics. A summary of open questions on these topics can be found in [11].
In this paper, we provide an answer to[11, Open Problem 2] in dimension $3$ and $4$. We study on locally finitely piecewise affine homeomorphisms $f:\Omega\to\Delta$, where $\Omega,\Delta\subset\mathbb{R}^{d}$, $d=3,4$ (for the definition see Definition 2.1). The key role these maps play in several areas make them natural objects of study as they bridge the gap between purely topological and differentiable structures. Our main result is the following:
Theorem A.
Let $d=3$ or $d=4$. Let $\Omega\subset\mathbb{R}^{d}$ be a domain and let $p,q\in[1,\infty)$. Let $f:\Omega\to\mathbb{R}^{d}$ be a locally finite piece-wise affine homeomorphism. Then for every $\varepsilon>0$ there exists a diffeomorphism $\tilde{f}$ satisfying
$\|\tilde{f}-f\|_{L^{\infty}(\Omega)}+\|\tilde{f}^{-1}-f^{-1}\|_{L^{\infty}(f(\Omega))}<\varepsilon$
and
$\|D\tilde{f}-Df\|_{L^{p}(\Omega)}+\|D\tilde{f}^{-1}-Df^{-1}\|_{L^{q}(f(\Omega))}<\varepsilon.$
Recently a body of research has accrued around the approximation of homeomorphisms by diffeomorphisms in Sobolev spaces. Let us note some of the most important results. The celebrated breakthrough results in the area, which stimulated much interest in the subject, were given by Iwaniec, Kovalev and Onninen in [14] and [15], where they found diffeomorphic approximations to any planar homeomorphism $f\in W^{1,p}(\Omega,\mathbb{R}^{2})$, for any $1<p<\infty$ in the $W^{1,p}$ norm. The remaining open case (in the plane), $p=1$, was solved by Hencl and Pratelli [12] using a different method, first generating piecewise affine approximations and then smoothing them.
All previous results which deal with the approximation of homeomorpshisms by diffeomorphisms are planar in nature. To this point there are no results allowing us to approximate Sobolev homeomorphisms by diffeomorphisms in the Sobolev space in higher dimension (to the knowledge of the authors). The construction in [14] yields a diffeomorphism whereas the construction in [12] generates a piecewise affine homeomorphism. Therefore, we quickly arrive at the question of whether approximation by piecewise affine homeomorphisms or by diffeomorphisms is equivalent. One implication of this question was shown in [17]. Here they show that any diffeomorphism can be triangulated for any dimension $d\geq 1$. The maps generated by their process remain close to the original diffeomorphism in $W^{1,\infty}$ and the same holds for the inverse. On the other hand, it is impossible for a diffeomorphism to be arbitrarily close to a piecewise affine homeomorphism in $W^{1,\infty}$ since the derivative must be continuous and this adds some extra difficulty to the opposite implication.
Known results in this direction, so far, are as follows. In [19] it was proven that in the planar case one can approximate a Sobolev piecewise affine map by diffeomorphisms in the Sobolev space. A further planar result working also with the second derivatives (of the map but not its inverse) has been achieved in [4]. It was proven in [6] that one can approximate any Sobolev piecewise affine homeomorphism by smooth homeomorphisms (this does not require that the approximations have smooth inverse) from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$ for any $d$.
One disadvantage of the previously mentioned result is that it is known that the concepts like connectedness and closure are quite different when we require only smooth homeomorphisms as opposed to diffeomorphisms, we refer the reader to [23]. Moreover, the Theorem in [23] implies that diffeomorphisms are connected in pointwise $\mathcal{C}^{k}$ topology, but this is not true in $\mathcal{C}^{k}$ uniform topology for all dimensions as can be observed from [9, Section II] and [13, Table 1 and 2] and connectedness of diffeomorphisms is a critical part of our construction. With respect to the technique in [12], the question of the equivalence of approximation by piecewise affine homeomorphisms and diffeomorphisms is very interesting in higher dimension and our result covers precisely this gap in dimension 3 and 4.
Theorem A can be generalized to the more general case in the setting of rearrangement invariant function spaces. For a definition of a rearrangement invariant function space $X$ with absolutely continuous norm, see section 2.1.
Theorem B.
Let $\Omega\subset\mathbb{R}^{d}$ be a domain. Let $f:\Omega\to\mathbb{R}^{d}$ be a locally finite piece-wise affine homeomorphism. Let $X(\Omega),Y(f(\Omega))$ be a pair of rearrangement invariant Banach function spaces with $\varphi_{X}(t),\varphi_{Y}(t)\to 0$ as $t\to 0$. Then for every $\varepsilon>0$ there exists a diffeomorphism $\tilde{f}\in\mathcal{C}^{1}(\Omega,\mathbb{R}^{n})$ with
$\|\tilde{f}-f\|_{L^{\infty}(\Omega)}+\|\tilde{f}^{-1}-f^{-1}\|_{L^{\infty}(f(\Omega))}<\varepsilon$
and
$\|D\tilde{f}-Df\|_{X(\Omega)}+\|D\tilde{f}^{-1}-Df^{-1}\|_{Y(f(\Omega))}<\varepsilon.$
In section 2 we introduce some notion and notation we need through the paper. Later, in section 3 we give a full detailed proof of the claim in dimension three. Finally we explain the differences for dimension four and comment on why we cannot directly generalize our approach to higher dimension.
2. Preliminaries
In this Section we introduce some notions and notations that we will use repeatedly.
Let $\eta\in\mathcal{C}^{\infty}(\mathbb{R})$ with
(2.1) $\eta(t)=0\text{ for all }t\leq 0,\eta(t)=1\text{ for all }t\geq 1\text{ and }0\leq\eta^{\prime}\leq 2.$
Definition 2.1.
Let $A_{1},A_{2},\dots A_{n+1}\in\mathbb{R}^{n}$ be points so that the vectors $A_{2}-A_{1}$, $A_{3}-A_{1},\dots A_{n+1}-A_{1}$ are linearly independent. Then we call $T=\operatorname{co}\{A_{1},A_{2},\dots,A_{n+1}\}$, the convex hull of the points $A_{1},A_{2},\dots,A_{n+1}$ a simplex (or more specifically an $n$-simplex).
Let $\{B_{1},B_{2},\dots,B_{k+1}\}\subset\{A_{1},A_{2},\dots A_{n+1}\}$ then we refer to $F=\operatorname{co}\{B_{1},B_{2},\dots,B_{k+1}\}$ as a $k$-subsimplex of $T$. Given $F$, a $k$-subsimplex of $T$ we denote by $N_{F}$ the space of vectors perpendicular to the vectors $B_{2}-B_{1}$, $B_{3}-B_{1},\dots,B_{k+1}-B_{1}$.
Let $\Omega\subset\mathbb{R}^{n}$, by the term simplicial complex of $\Omega$ we refer to a countable (possibly finite) set $\mathcal{T}=\{T_{1},T_{2},\dots\}$ of simplices such that every $k$-subsimplex of a simplex from $\mathcal{T}$ is also in $\mathcal{T}$, every non-empty intersection of any two simplices in $\mathcal{T}$ is a $k$-subsimplex of both of them and $\Omega=\bigcup_{i}T_{i}$.
We call a simplicial complex locally finite if for any $x\in\Omega$ there exists an $r>0$ such that only a finite number of simplices $T_{i}$ intersect $B(x,r)$.
We refer to $f$ as a piecewise affine map on $\Omega$ if there exists a $\mathcal{T}$ simplicial complex on $\Omega$ such that the restriction of $f$ on each $T_{i}\in\mathcal{T}$ is equal to some affine map $L_{i}(x)=M_{i}x+c_{i}$. Further we say that $f$ is a locally finite piecewise affine homeomorphism, if $\mathcal{T}$ is locally finite.
For a vector $v\in\mathbb{R}^{3}$ by $\partial_{v}f$ we denote the directional derivative of $f$ in the direction $v$. By $\pi_{3}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ we denote the projection
$\pi_{3}(x,y,z)=(x,y,0)$
and similarly by $\tilde{\pi}_{3}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ we denote the projection
$\tilde{\pi}_{3}(x,y,z)=(x,y).$
The following theorem follows from [24, Theorem 6].
Theorem 2.2.
For every $\alpha_{0}$, a sense-preserving $\mathcal{C}^{\infty}$ diffeomorphism of $S^{2}=\partial B_{\mathbb{R}^{3}}(0,1)$ onto itself, there exists an $\alpha\in\mathcal{C}^{\infty}([0,1]\times S^{2},S^{2})$ such that $\alpha(t,\cdot)$ is a sense-preserving diffeomorphism of $S^{2}$ for each $t$, $\alpha(0,\cdot)=\operatorname{id}(\cdot)$ and $\alpha(1,\cdot)=\alpha_{0}(\cdot)$.
Corollary 2.3.
Let $\tilde{f}$ be a $C^{\infty}$ sense-preserving diffeomorphism of $S^{2}=\partial B_{\mathbb{R}^{3}}(0,1)$ onto itself. There exists $g$, a $C^{\infty}$ diffeomorphism of $\overline{B_{\mathbb{R}^{3}}(0,1)}$ onto $\overline{B_{\mathbb{R}^{3}}(0,1)}$ with $g_{\rceil\partial B_{\mathbb{R}^{3}}(0,1)}=\tilde{f}$.
Proof.
By Theorem 2.2 we have an isotopy $h:\partial B(0,1)\times[0,1]\to\partial B(0,1)$ such that $h(x,t)$, $t\in[0,1]$, $h(x,1)=\tilde{f}(x)$ and $h(x,0)=x$ for all $x\in\partial B(0,1)$. In fact may assume that $h(x,t)=\tilde{f}(x)$ for $t\in[\tfrac{2}{3},1]$ and $h(x,t)=x$ for $t\in[0,\tfrac{1}{3}$]. Then we define $g:\overline{B(0,1)}\to\overline{B(0,1)}$ as
$g(x)=\begin{cases}|x|h\Big{(}\frac{x}{|x|},|x|\Big{)}\quad&|x|>0\\ 0&x=0.\end{cases}$
Then $g$ is a diffeomorphism and $g=\tilde{f}$ on $\partial B(0,1)$. ∎
The following lemma is well-known, we include it for simple reference.
Lemma 2.4.
Let $d\geq 1$, let $\Omega,\Delta\subset\mathbb{R}^{d}$ be bounded domains and let $\mu:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be continuous and a local diffeomorphism of $\Omega\subset\mathbb{R}^{d}$. Further we assume that there is a neighborhood $O$ of $\partial\Omega$ in $\overline{\Omega}$ and a neighborhood $D$ of $\partial\Delta$ in $\overline{\Delta}$ such that $\mu$ is a homeomorphism $O$ onto $D$. Then $\mu$ is a diffeomorphism of $\Omega$ onto $\Delta$.
Proof.
It is a well known fact from degree theory that $\deg(\mu,\cdot,\Omega)$ is constant on components of $\mathbb{R}^{n}\setminus\mu(\partial\Omega)$. Also it is well known that the degree of a homeomorphism is ±$\pm 1$. Since $\mu$ is equal to a homeomorphism on $D$, we have that $\deg(\mu,\cdot,\Omega)=\pm 1$ on $\Delta$ and is constant. Since $\mu$ is a local diffeomorphism, the sign of $\det D\mu$ is constant (and equal to $\deg(\mu,y,\Omega)$ for every $y\in\Delta$) on $\Omega$. Since every point $y\in\Delta$ is a regular value of $\mu$, there is at most one point $x\in\Omega$ such that $\mu(x)=y$. Similarly for all $y\in\mathbb{R}^{d}\setminus\overline{\Omega}$ it holds that $\mu^{-1}(\{y\})=\emptyset$. Thus $\mu$ is injective on $\Omega$ and $\mu(\Omega)=\Delta$. It follows that $\mu$ is a diffeomorphism of $\Omega$ onto $\Delta$. ∎
Let $\mu:S^{d-1}\to S^{d-1}$ be a smooth mapping. We call the derivative of $\mu$, the linear map $d\mu_{x}:T_{x}S^{d-1}\to T_{\mu(x)}S^{d-1}$ defined in the usual way. Let $y\in S^{d-1}$ be such that the determinant of $d\mu_{x}$ is non-zero for all $x\in\mu^{-1}(y)$ (which is true for almost all points $y\in\mu(S^{d-1})$) then we define
$\deg(\mu,y):=\sum_{x\in\mu^{-1}(y)}\operatorname{sgn}\det d\mu_{x}.$
For more details we refer the reader to [18, Chapter 5].
Theorem 2.5.
Let $d\geq 2$ and $\mu:S^{d-1}\to S^{d-1}$ be a local diffeomorphism and let there be a point $Q\in S^{d-1}$ such that $\deg(\mu,Q)=1$, then $\mu$ is a diffeomorphism.
Proof.
Since $\mu$ is a local diffeomorphism, it is a proper map over $S^{d-1}$. By [8, Proposition 5.12], it holds that $\deg(\mu,y)=\deg(\mu,Q)=1$ for all $y$ in $S^{d-1}$. Since $\mu$ is a local diffeomorphism, it follows (as in the proof of Lemma 2.4), that for each $y\in S^{d-1}$ there is exactly one preimage $x\in S^{d-1}$. Thus, $\mu$ is a diffeomorphism of $S^{d-1}$ onto $S^{d-1}$. ∎
2.1. Rearrangement invariant function spaces
Let $E$ be a (non-negligible) Lebesgue measurable subset of $\mathbb{R}^{d}$ of finite measure. We denote by $\mathcal{L}^{d}(E)$ its Lebesgue measure. We set
$L^{0}(E)=\left\{f:f\text{ is measurable function on }E\text{ with values in }[-\infty,+\infty]\right\}$
and
$L^{0}_{+}(E)=\left\{f\in L^{0}(E):f\geq 0\right\}.$
We identify functions $f_{1},f_{2}$ for which $\mathcal{L}^{d}(\{f_{1}\neq f_{2}\})=0$ in the space $L^{0}$. The non-increasing rearrangement $f^{*}\colon[0,+\infty]\rightarrow[0,+\infty]$ of a function $f\in L^{0}(E)$ is defined by
$f^{*}(s)=\inf\left\{t\geq 0:\mathcal{L}^{d}\big{(}\left\{x\in E:|f(x)|>t\right\}\big{)}\leq s\right\}\qquad s\in[0,+\infty)$
and the Hardy-Littlewood maximal function $f^{**}:(0,\mathcal{L}^{d}(E))\to[0,+\infty)$ is given by
[…]
Definition 2.6.
Let $E$ be a Lebesgue measurable subset of $\mathbb{R}^{d}$ and let $\|\cdot\|_{X(E)}:L^{0}(E)\rightarrow[0,+\infty]$ be a functional. Consider the following properties
  1. (P1)
    $\|\cdot\|_{X(E)}$ is a norm on $L^{0}(E)$.
  2. (P2)
    For all $f,g\in L^{0}_{+}(E)$ the inequality $f(x)\leq g(x)$ for a.e. $x$ in $E$ implies $\|f\|_{X(E)}\leq\|g\|_{X(E)}$.
  3. (P3)
    $\sup\limits_{k}\|f_{k}\|_{X(E)}=\|f\|_{X(E)}$ if $0\leq f_{k}(x)\nearrow f(x)$ for a.e. $x$ in E.
  4. (P4)
    Let $G\subset E$ be a set of a finite measure. Then
    $\|\chi_{G}\|_{X(E)}<\infty.$
  5. (P5)
    Let $G\subset E$ be a set of a finite measure. Then there exists a constant $C_{G}$ depending only on the choice of the set $G$ for which
    $\|f\chi_{G}\|_{L^{1}(E)}\leq C_{G}\|f\chi_{G}\|_{X(E)}.$
    for all $f\in L^{0}(E)$.
  6. (P6)
[…]
If $\|\cdot\|_{X}$ enjoys the properties (P1)-(P5) we call it a Banach function norm. If it also enjoys (P6) we call it a rearrangement invariant Banach function norm. Let $\|\cdot\|_{X}$ be (rearrangement-invariant) Banach function norm then we call set
$X(E):=\{f\in L^{0}(E):\|f\|_{X(E)}<\infty\}$
endowed with the norm $\|\cdot\|_{X(E)}$ a (rearrangement-invariant) Banach function space.
Given a r.i. Banach function space $X(E)$ and $0\leq s<\mathcal{L}^{d}(E)$ one may define the fundamental function of $X(E)$ by
(2.2) $\varphi_{X(E)}(s):=\|\chi_{G}\|_{X(E)}$
where $G\subset E$ is an arbitrary subset of $E$ of measure $s$.
Let $X(E)$ be a Banach function space. We say that $X(E)$ has locally absolutely continuous norm if for any finite measure set $M\subset E$ and any function $f\in X(E)$ one has
$M\supset M_{n}\textup{such that }\mathcal{L}^{d}(M_{n})\rightarrow 0\quad\textup{implies}\quad\|f\chi_{M_{n}}\|_{X}\rightarrow 0.$
The following lemma is from [3].
Lemma 2.7.
Let $G\subset\mathbb{R}^{n}$ be a set of finite measure and $X(G)$ be a rearrangement invariant Banach function space satisfying
(2.3) ${\lim_{t\rightarrow 0+}}\varphi_{X}(t)=0.$
Then for every $M>0$ and $\tilde{\varepsilon}>0$ there exists a $\tilde{\delta}>0$ such that for all $u\in X(G)$ with $\|u\|_{L^{\infty}(G)}\leq M$ and $\|u\|_{L^{1}(G)}<\tilde{\delta}\mathcal{L}^{d}(G)$ one has $\|u\|_{X(G)}<\tilde{\varepsilon}\mathcal{L}^{d}(G)$.
3. The construction in 3D
In this Section we present a detailed proof of our result in dimension $d=3$, we start with the following Lemma that it introduces a map that we will use repeatedly through the paper.
Lemma 3.1.
Let $A_{1},A_{2}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ be affine be such that the mapping
$f(x)=\begin{cases}A_{1}x\quad&\text{for }x_{1}\leq 0\\ A_{2}x\quad&\text{for }x_{1}\geq 0.\end{cases}$
is a sense-preserving homeomorphism. Further, let $(D_{1}A_{2})_{1}\geq(D_{1}A_{1})_{1}>0$ and $w\in\mathcal{C}^{\infty}(\mathbb{R}^{2})$ be a positive function. Then there exists an $\sigma>0$ such that whenever $|Dw(y,z)|\leq\sigma$ the map
(3.1) $g_{w}(x)=\Big{(}1-\eta\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}\Big{)}A_{1}x+\eta\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}A_{2}x.$
is a diffeomorphism $g_{w}:\mathbb{R}\to\mathbb{R}^{3}$ satisfying
(3.2) $g_{w}(x_{1},x_{2},x_{3})=f(x_{1},x_{2},x_{3})$
on the set
$\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3};x_{1}\leq 0\}\cup\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3};x_{1}\geq w(x_{2},x_{3})\}$
[…]
Proof.
Since $A_{1}$, $A_{2}$, $w$ and $\eta$ are all in $\mathcal{C}^{\infty}$ and $w>0$ everywhere we have that $g_{w}\in\mathcal{C}^{\infty}(\mathbb{R}^{3})$. Further from definition, it is immediately obvious that $g_{w}(x_{1},x_{2},x_{3})=f(x_{1},x_{2},x_{3})$ on the set $\{x_{1}\leq 0\}\cup\{x_{1}\geq w(x_{2},x_{3})\}$.
Note that, since $A_{1}=A_{2}$ on the plane $\{x\in\mathbb{R}^{3};x_{1}=0\}$, we have $D_{2}A_{1}=D_{2}A_{2}$ and $D_{3}A_{1}=D_{3}A_{2}$. On the other hand the definition of $g_{w}$ is independent of addition by a vector, i.e. let $v\in\mathbb{R}^{3}$ and let $\tilde{f}=f+v$, then $\tilde{g}_{w}=g_{w}+v$. The same holds regarding the action of linear isometries, $O(3)$. Therefore, without loss of generality we can assume that $f$ maps the plane $\{x_{1}=0\}$ onto itself. This implies that $(D_{2}Q_{1})_{1}=(D_{3}Q_{1})_{1}=(D_{2}Q_{2})_{1}=(D_{3}Q_{2})_{1}=0$. Further, we may assume that $f$ has positive Jacobian.
Firstly, we consider the simplified case where $w>0$ is an arbitrary constant. We prove that, in this case, our construction yields a diffeomorphism and from this we obtain the existence of $\sigma>0$ by the continuity of the Jacobian of $g_{w}$ with respect to $Dw$.
Let us now assume that $w>0$ is constant. From the above and the definition of $g_{w}$ we have that
(3.4) $D_{2}g_{w}=D_{2}f\qquad\text{and}\qquad D_{3}g_{w}=D_{3}f.$
Therefore, we have that $g_{w}$ maps all planes of type $\{x_{1}=c\}$ onto a plane $\{x_{1}=\tilde{c}(c)\}$ and this restriction is a regular two-dimensional affine map. In order to prove that $g_{w}$ is injective it suffices to prove that $[D_{1}g_{w}]^{1}>0$. We have that
(3.5) $[D_{1}g_{w}]^{1}=$ $\Big{(}1-\eta\Big{(}\frac{x_{1}}{w}\Big{)}\Big{)}[D_{1}A_{1}]^{1}+\eta\Big{(}\frac{x_{1}}{w}\Big{)}[D_{1}A_{2}]^{1}$
$\quad+\frac{1}{w}\eta^{\prime}\Big{(}\frac{x_{1}}{w}\Big{)}[A_{2}x_{1}-A_{1}x_{1}]^{1}$
$\geq$ $[D_{1}A_{1}]^{1}+\frac{1}{w}\eta^{\prime}\Big{(}\frac{x_{1}}{w}\Big{)}[A_{2}x_{1}-A_{1}x_{1}]^{1}$
$\geq$ $[D_{1}A_{1}]^{1}>0.$
Therefore, $J_{g_{w}}\geq[D_{1}A_{1}]^{1}J_{2}^{f}>0$, where, by $J_{2}^{f}$, we denoted the determinant of the $2\times 2$ matrix $[D_{i}f]^{j}$, $i,j=1,2$. Thus, we see that $g_{w}$ is a diffeomorphism in the case that $w$ is constant.
Let us now assume that $w$ is a smooth positive function and we find $\sigma$ such that $J_{g_{w}}\geq C>0$, whenever $|Dw|\leq\sigma$. We calculate $Dg_{w}$ as
(3.6) $D_{1}g_{w}(x)=\Big{(}1-$ $\eta\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}\Big{)}D_{1}A_{1}+\eta\Big{(}\frac{x}{w(x_{2},x_{3})}\Big{)}D_{1}A_{2}$
$+\frac{1}{w(x_{2},x_{3})}\eta^{\prime}\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}(A_{2}x-A_{1}x),$
and for $j=2,3$ we have
(3.7) $D_{j}g_{w}(x)$ $=\Big{(}1-\eta\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}\Big{)}D_{j}A_{1}+\eta\Big{(}\frac{x}{w(x_{2},x_{3})}\Big{)}D_{j}A_{2}$
$\quad+\frac{x_{1}}{w^{2}(x_{2},x_{3})}\eta^{\prime}\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}D_{j}w(x_{2},x_{3})(A_{2}x-A_{1}x)$
$=D_{j}f(x)-\frac{x_{1}}{w^{2}(x_{2},x_{3})}\eta^{\prime}\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}D_{j}w(x_{2},x_{3})(A_{2}x-A_{1}x).$
Note that there exists some $\Sigma>0$ such that
$\frac{x_{1}}{w^{2}(x_{2},x_{3})}\eta^{\prime}\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}(A_{2}x-A_{1}x)\leq\Sigma$
since, either $\eta^{\prime}\Big{(}\frac{x_{1}}{w(x_{2},x_{3})}\Big{)}=0$ or $x_{1}<w(x_{1},x_{2})$ and $|A_{2}x-A_{1}x|<|A_{2}-A_{1}|w(x_{2},x_{3})$. Together, this implies that
$J_{g_{w}}(x)\geq[D_{1}A_{2}-D_{1}A_{1}]^{1}J_{2}^{f}-C\|Df\|_{L^{\infty}}\big{(}\Sigma|Dw|C\|Df\|_{L^{\infty}}+\Sigma^{2}|Dw|^{2}\big{)}.$
Therefore, we find some $\sigma>0$ such that $J_{g_{w}}\geq\tfrac{1}{2}[D_{1}A_{2}-D_{1}A_{1}]^{1}J_{2}^{f}$.
Let us now prove that $J_{g_{w}}\geq C>0$ implies that $g_{w}$ is a diffeomorphism. It suffices to show that the restriction of $g_{w}$ to $B(0,R)$ is a homeomorphism for every $R$. Let $R>0$ be fixed. We find a $\tilde{w}\in\mathcal{C}^{\infty}(\mathbb{R}^{2})$ with $|D\tilde{w}|\leq\sigma$, $\tilde{w}(y,z)=w(y,z)$ for $(y,z)\in B_{\mathbb{R}^{2}}(0,R)$ and $\tilde{w}$ is constant outside $B(0,4R)$. Then $g_{\tilde{w}}$ equals a (sense-preserving) homeomorphism on the boundary of any cube, which contains $B(0,4R)$. By Theorem 2.4 $g_{\tilde{w}}$ is a homeomorphism. But since $g_{w}=g_{\tilde{w}}$ on $B(0,R)$ we have that $g_{w}$ is a homeomorphism on $B(0,R)$.
[…]
In the following lemma we use a rotated and possibly reflected version of the previous lemma. Let us be describe this in greater detail. Let $\theta_{i}\in[-\pi,\pi)$ we divide the $\mathbb{R}^{3}$ by a plane generated by $(0,0,1)$ and $(\cos\theta_{i},\sin\theta_{i},0)$ after a rotation of $-\theta_{i}$ in the $x,y$-plane this becomes the plane $\{x_{1}=0\}$. We assume that we have two affine mappings related to this plane $A_{i}$ and $A_{i+1}$ such that
(3.8) $f_{i}(x)=\begin{cases}A_{i}x\quad&\text{for }\langle(x_{1},x_{2}),(\cos\theta_{i},\sin\theta_{i})\rangle\leq 0\\ A_{i+1}x\quad&\text{for }\langle(x_{1},x_{2}),(\cos\theta_{i},\sin\theta_{i})\rangle\geq 0.\end{cases}$
is a sense-preserving homeomorphism. Further we assume that we have a smooth positive map $w_{i}$ on the plane containing the set $\{\arg(x_{1},x_{2})=\theta_{i}\}$. Call $\tilde{A}_{i}$ the composition of the rotation by $\theta_{i}$ in the $x,y$ plane composed with $A_{i}$ and $\tilde{A}_{i+1}$ similarly. Also, we define $\tilde{w}_{i}$ as the composition of the rotation and $w_{i}$.
If the corresponding mapping $\tilde{f}$ satisfies $[D_{1}\tilde{A}_{2}]^{1}\geq[D_{1}\tilde{A}_{1}]^{1}>0$ then we can directly construct a map $\tilde{g}_{\tilde{w}_{i}}$ exactly as in the claim of the previous lemma. In the case that $0<[D_{1}A_{2}]^{1}<[D_{1}A_{1}]^{1}$ it is sufficient a reflection in $\{x_{1}=0\}$ in the image and in the pre-image returns us to the previous case. The other cases can be dealt with a single reflection. After defining $\tilde{g}_{\tilde{w}_{i}}$
(3.9) we define gi as the reflection and rotation of $\text{we define $g_{i}$ as the reflection and rotation of $\tilde{g}_{\tilde{w}_{i}}$}.$
$B_{\mathbb{R}^{2}}(0,r)$
Figure 1. The red areas are the sets where $\operatorname{dist}(x,\mathbb{R}(\cos\theta_{i},\sin\theta_{i}))<w_{i}$. If $w_{i}$ is small compared to $r$ then these sets are disjoint outside $B(0,r)$. Each $g_{i}$ is a diffeomorphism of $\mathbb{R}^{3}$ as is the map $f$ on $\{\operatorname{dist}((x_{1},x_{2}),\mathbb{R}(\cos\theta_{i},\sin\theta_{i}))<w_{i}\}$. Since $g_{i}=g_{i+1}=f$ on the set $\arg(x_{1},x_{2})\in(\theta_{i},\theta_{i+1})$ outside $\{\operatorname{dist}(x,\mathbb{R}(\cos\theta_{i},\sin\theta_{i}))<w_{i}\}$, we get that the map $g$ defined in (3.11) is a diffeomorphism.
If we assume that $w_{i}$ is small compared to the distance to the line ×{(0,$\mathbb{R}\times\{(0,0)\}$, i.e.
(3.10) $\arctan\Big{(}\frac{w_{i}(x_{1},x_{2})}{\sqrt{x_{1}^{2}+x_{2}^{2}}}\Big{)}<\frac{\min\{|\theta_{i}-\theta_{j}+k\pi|;1\leq i<j\leq m;\ k\in\mathbb{Z}\}\cup\{\tfrac{\pi}{8}\}}{8},$
for all $x_{1}^{2}+x_{2}^{2}\geq r^{2}$ then we have that the mapping defined as
(3.11) $g(x):=g_{i}(x)\quad\arg(x_{1},x_{2})\in[\tfrac{1}{2}\theta_{i-1}+\tfrac{1}{2}\theta_{i},\tfrac{1}{2}\theta_{i}+\tfrac{1}{2}\theta_{i+1}]+2k\pi$
for $k\in\mathbb{Z}$, is a diffeomorphism on $\{x\in\mathbb{R}^{3};x_{1}^{2}+x_{2}^{2}\geq r^{2}\}$. This is because $g_{i}=g_{i+1}=f$ on the set
$\{\arg(x_{1},x_{2})\in(\theta_{i},\theta_{i+1})\}\setminus\{\operatorname{dist}((x_{1},x_{2}),\mathbb{R}(\cos\theta_{i},\sin\theta_{i}))<w_{i}\},$
and the sets
$\{x\in\mathbb{R}^{3};\operatorname{dist}(x,\mathbb{R}(\cos\theta_{i},\sin\theta_{i}))<w_{i}\}\setminus B(0,r)$
are pairwise disjoint (since the condition (3.10) ensures that the sets $\{x\in\mathbb{R}^{2}\setminus B(0,r);\operatorname{dist}(x,\mathbb{R}(\cos\theta_{i},\sin\theta_{i}))<w_{i}\}$ lie in disjoint cones around the rays $\mathbb{R}(\cos\theta_{i},\sin\theta_{i})$). We refer the reader to Figure 1. In fact, if the estimate in (3.10) holds for $x_{1}^{2}+x_{2}^{2}\geq r^{2}/16$ then the sets $\{\operatorname{dist}(x,\mathbb{R}(\cos\theta_{i},\sin\theta_{i}))<w_{i}\}\setminus B(0,r/4)$ are pairwise disjoint. In this case we have that the mapping
(3.12) $\tilde{g}(x):=g_{i}(x)\quad\arg(x_{1},x_{2})\in[\tfrac{1}{2}\theta_{i-1}+\tfrac{1}{2}\theta_{i},\tfrac{1}{2}\theta_{i}+\tfrac{1}{2}\theta_{i+1}]+2k\pi$
is a diffeomorphism on $\{x\in\mathbb{R}^{3};x_{1}^{2}+x_{2}^{2}\geq r^{2}/4\}$.
Lemma 3.2.
Let $m\in\mathbb{N}$, let $-\pi\leq\theta_{1}<\theta_{2}<\dots<\theta_{m}<\pi$, and $A_{1},A_{2}\dots A_{m}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ be linear, such that
$f(x)=A_{i}x\quad\arg(x_{1},x_{2})\in\bigcup_{k\in\mathbb{Z}}[\theta_{i}+2k\pi,\theta_{i+1}+2k\pi].$
[…]
Proof.
Step 1. Setup.
We use the function $\eta$ from (2.1). After a rotation in the image (and possibly a reflection) we may assume that $D_{3}f=(0,0,\lambda)$ for some $\lambda>0$, where $f$ is the piecewise affine map from (3.8) used to construct $g$. Recall that we denote $\pi_{3}(x_{1},x_{2},x_{3})=(x_{1},x_{2},0)$ and $\tilde{\pi}_{3}(x_{1},x_{2},x_{3})=(x_{1},x_{2})$.
Let us outline the rough idea of how we proceed in the following. We start by defining $g$ on $S:=\{(0,0)\}\times\mathbb{R}$ as $f(0,0,x_{3})=x_{3}D_{3}f$. For brevity of notation we write $\Phi(t,\theta,x_{3})=(t\cos\theta,t\sin\theta,x_{3})$. In the following we define $g$ on concentric cylindrical annuli around $S$ called $P_{1},P_{2},P_{3}$. On the outer annulus, $P_{1}$ we define $g$ as a convex combination which “flattens” the images annuli where $x_{3}=c$ so that their image lies on the planes perpendicular to the vector $(0,0,1)$ (i.e. the planes are perpendicular to $D_{3}f$). The image of the outer boundary circle of $P_{2}\cap\{x_{3}=c\}$ is a smooth star-shaped domain. On $P_{2}$ we do a radial squeezing so that the image of the inner boundary circle of $P_{2}\cap\{x_{3}=c\}$ is a circle. On $P_{3}$ we use an isotopy between a diffeomorphism of $S^{1}$ and a rotation. Then on the remaining cylinder inside we define $g$ as the rotation from the previous step. We refer the reader to Figure 2 for an illustration of the idea.
Figure 2. In the annulus $P_{1}$ we shift in the $e_{3}$ direction to so that hyperplanes disks inside $P_{1}$ are mapped onto hyperplanes. In $P_{2}$ we squeeze so that the image of a circle is a circle. In $P_{3}$ we ‘untwist’ to achieve a rotation in the disks inside $P_{3}$ we extend as the rotation.
Step 2. Flatten images of hyperplanes pieces onto hyperplane pieces.
Let us start with the simplified case where $w_{i}$ are constant and satisfy (3.10). On $P_{1}:=\{(t\cos\varphi,t\sin\varphi,x_{3}):x_{3}\in\mathbb{R},t\in[\tfrac{4}{5}r,r]\varphi\in\mathbb{R}\}$ we define
$g(\Phi(t,\varphi,x_{3}))=$ $\Big{[}1-\eta\Big{(}\frac{5t-4r}{r}\Big{)}\Big{]}\Big{(}(0,0,\lambda x_{3})+\pi_{3}\circ\tilde{g}\big{(}\Phi(t,\theta,x_{3})\big{)}\Big{)}$
$+\eta\Big{(}\frac{5t-4r}{r}\Big{)}\tilde{g}\big{(}\Phi(t,\theta,x_{3})\big{)}.$
Now we prove that the map $g$ is injective on $P_{1}$. Because $D_{3}f=(0,0,\lambda)$ on $\mathbb{R}^{3}$, it holds that $D_{3}\tilde{g}=(0,0,\lambda)$ on $\mathbb{R}^{3}\setminus[B_{\mathbb{R}^{2}}(0,r/4)\times\mathbb{R}]$ (see (3.4), we used the fact that $w_{i}$ are all constant) where $\tilde{g}$ is the map from (3.12). Then we easily also observe that $D_{3}g=(0,0,\lambda)$ everywhere on $P_{1}$. Since we may assume that $J_{\tilde{g}}>0$, we also have that
$\det\left(\begin{matrix}D_{1}{g}_{1}(x)&D_{2}{g}_{1}(x)\\ D_{1}{g}_{2}(x)&D_{2}{g}_{2}(x)\end{matrix}\right)=\det\left(\begin{matrix}D_{1}\tilde{g}_{1}(x)&D_{2}\tilde{g}_{1}(x)\\ D_{1}\tilde{g}_{2}(x)&D_{2}\tilde{g}_{2}(x)\end{matrix}\right)\geq C>0$
for some $C$ and all $x\in\mathbb{R}^{3}\setminus[B_{\mathbb{R}^{2}}(0,r/4)\times\mathbb{R}]$. Then we may conclude that $J_{g}>0$ on $P_{1}$. It is not hard to observe that $g$ is injective on $\partial P_{1}$. In fact it suffices to prove that $g$ is injective on the the circle $\partial B_{\mathbb{R}^{2}}(0,3r/4)\times\{0\}$ which is easily observed from (3.5). From the definition of $g$ we immediately see that $g$ is smooth on $P_{1}$ and has Jacobian bounded from below above zero and so is a diffeomorphism.
Step 3. Squeeze the smooth star-shaped domain radially into a small circle.
Call $P_{2}:=\Phi([\tfrac{3}{5}r,\tfrac{4}{5}r],\mathbb{R},\mathbb{R})$. We find a number $\rho>0$ so small that
$B_{\mathbb{R}^{2}}(0,\rho)\subset\mathbb{R}^{2}\setminus\tilde{\pi}_{3}\circ\tilde{g}(\{x_{1}^{2}+x_{2}^{2}>r^{2}/4\}).$
Then we define
$g(\Phi(t,\theta,x_{3}))=$ $(0,0,\lambda x_{3})+\eta\Big{(}\frac{5t-3r}{r}\Big{)}\pi_{3}\circ\tilde{g}\big{(}\Phi(t,\theta,x_{3})\big{)}$
$+\Big{(}1-\eta\Big{(}\frac{5t-3r}{r}\Big{)}\Big{)}t\rho\frac{\pi_{3}\circ\tilde{g}\big{(}\Phi(\tfrac{3}{5}r,\theta,x_{3})\big{)}}{|\pi_{3}\circ\tilde{g}\big{(}\Phi(\tfrac{3}{5}r,\theta,x_{3})\big{)}|}.$
The question of whether $g$ is injective is clearly only a 2-dimensional one, since disjoint planes have disjoint images i.e.
$g\big{(}\Phi([\tfrac{3}{5}r,\tfrac{4}{5}r],\mathbb{R},a)\big{)}\cap g\big{(}\Phi([\tfrac{3}{5}r,\tfrac{4}{5}r],\mathbb{R},b)\big{)}=\emptyset\quad\text{ if }a\neq b.$
It is easily observed that radial segments in $P_{2}$ are mapped to radial segments in the image with derivative bounded from below by $C(f)$ thanks to the choice of $\rho$. As stated in the previous step, (3.5) allows us to observe that $\partial_{\theta}g(\Phi(t,\theta,x_{3}))$ has a component in a direction perpendicular to $g(x)/|g(x)|$ bounded away from zero, i.e. also greater than $\rho C_{f}$. Therefore, we observe that
$\det\left(\begin{matrix}D_{1}{g}_{1}(x)&D_{2}{g}_{1}(x)\\ D_{1}{g}_{2}(x)&D_{2}{g}_{2}(x)\end{matrix}\right)\geq C>0$
on $P_{2}$, where $C$ depends on $f$ but not on values of $w$ or $r$.
Step 4. ‘Untwist’ the circle until it becomes a rotation and extend as the rotation close to the segment $S$.
It is well known that for any sense preserving diffeomorphism $h$ of the circle $S^{1}$ onto $S^{1}$ there exists an isotopy $\tilde{\alpha}:S^{1}\times[0,1]\to S^{1}$ with $\tilde{\alpha}(\zeta,1)=h(\zeta)$ and $\tilde{\alpha}(\zeta,0)=\zeta$ and that
(3.14) $C^{-1}\inf_{\zeta\in S^{1}}|h^{\prime}(\zeta)|\leq|\partial_{1}\tilde{\alpha}(\tilde{\zeta},t)|\leq C\sup_{\zeta\in S^{1}}|h^{\prime}(\zeta)|$
for all $t\in[0,1]$ and $\tilde{\zeta}\in S^{1}$. Further
[…]
for all $t\in[0,1]$ and $\tilde{\zeta}\in S^{1}$. A proof and explicit construction of this can be found in [19, Section 3.1] but it is not hard to construct the map by hand using a smooth convex combination in polar coordinates. For us
$h(\cos\theta,\sin\theta)=\frac{\tilde{\pi}_{3}\circ\tilde{g}\big{(}\Phi(\tfrac{3}{5}r,\theta,x_{3})\big{)}}{|\tilde{\pi}_{3}\circ\tilde{g}\big{(}\Phi(\tfrac{3}{5}r,\theta,x_{3})\big{)}|}=\frac{\tilde{\pi}_{3}\circ g(\Phi(\tfrac{3}{5}r,\theta,x_{3}))}{|\tilde{\pi}_{3}\circ g(\Phi(\tfrac{3}{5}r,\theta,x_{3}))|}$
and so let $\alpha$ be the corresponding isotopy.
Call $P_{3}:=\Phi([\tfrac{2}{5}r,\tfrac{3}{5}r],\mathbb{R},\mathbb{R})$. We define
$g(\Phi(t,\theta,x_{3}))=$ $(0,0,\lambda x_{3})+t\rho\alpha\Big{(}\theta,\frac{5t-2r}{r}\Big{)}.$
Since $\alpha$ is a diffeomorphism, we have that $g$ is injective on $P_{3}$. Thus, we observe that $g$ is injective on $P_{1}\cup P_{2}\cup P_{3}$. Further, because $\alpha$ is smooth, $g$ is smooth and can be smoothly extended as
$g(\Phi(t,\theta,x_{3}))=$ $(0,0,\lambda x_{3})+t\rho\alpha(\theta,0)$
for $t\in[0.\tfrac{r}{4}]$. Notice that this agrees with our initial definition of $g$ on $S$.
By considering the definition of $g$ on each part $P_{1},P_{2},P_{3}$ separately, we observe that
(3.16) $D_{3}$ $g(\Phi(x_{1},t,\varphi))=(0,0,\lambda)\leq\|Df\|_{L^{\infty}},$
$|\partial_{t}$ $g(\Phi(t,\theta,x_{3}))|\leq C\frac{1}{r}\sup_{\theta\in\mathbb{R}}|\pi_{3}\circ\tilde{g}\big{(}\Phi(r,\theta,0)\big{)}|$
$\ \ +C\frac{1}{r}\sup_{\Phi(t,\theta,0)\in P_{1}}\Big{|}\Big{[}\tilde{g}\big{(}\Phi(t,\theta,0)\big{)}-\pi_{3}\circ\tilde{g}\big{(}\Phi(t,\theta,0)\big{)}\Big{]}\Big{|}$
[…]
$\qquad\qquad\quad\ \ \leq C\|Df\|_{L^{\infty}},\text{ and}$
$|\partial_{v}$ $g(\Phi(t,\theta,x_{3}))|\leq C\|Df\|_{L^{\infty}}$
where by $v=v(\theta)$ we denote the vector $(\sin\theta,-\cos\theta)$. These estimates suffice to give (3.13) since we the ratio $\rho/r$ is a fixed constant depending on $f$. We have proved also that $J_{g}\geq C>0$ implying that, thanks to Theorem 2.4, $g$ is a diffeomorphism since we know that $g$ is injective far away from $S$ by Lemma 3.1.
Step 5. The construction works also for non-constant $w_{i}$.
Let us now assume that the functions $w_{i}$ are not constant. The derivative $Dg$ depends continuously on $Dw_{i}$ and so we find a $\tilde{\sigma}>0$ so small that the Jacobian of $g$ remains bounded away from 0 as soon as $|Dw|<\tilde{\sigma}$. The conclusion now mirrors that above for constant $w_{i}$. ∎
Let us now prove that the construction from Lemma 3.2 can be modified to allow us to modify the parameter $r$ with respect to $x_{3}$ while still giving a diffeomorphic extension and the same estimates on $\|Dg\|_{L^{\infty}}$.
Corollary 3.3.
Let $f$ be the piecewise affine map from (3.8). There exists a $\xi>0$ such that the following holds. Let $r\in\mathcal{C}^{\infty}(\mathbb{R})$ be a positive function such that $|r^{\prime}|\leq\xi$ and let $w_{i}\in\mathcal{C}^{\infty}\{x\in\mathbb{R}^{3};\arg(x_{1},x_{2})=\theta_{i}\}$ satisfy $\|Dw_{i}\|_{L^{\infty}}\leq\min\{\sigma_{i},\tilde{\sigma}\}$ (the numbers from Lemma 3.1 and Lemma 3.2), let also $w_{i}$ satisfy (3.10) for all $(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$ such that $x_{1}^{2}+x_{2}^{2}\leq r^{2}(x_{3})/16$. Then the map $\tilde{g}$ from (3.12) is a diffeomorphism of $\mathbb{R}^{3}\setminus\{x\in\mathbb{R}^{3};x^{2}_{1}+x^{2}_{2}\leq r^{2}(x_{3})/4\}$ and by construction in Lemma 3.2 we get a diffeomorphism equaling $\tilde{g}$ on $\mathbb{R}^{3}\setminus\{x\in\mathbb{R}^{3};x^{2}_{1}+x^{2}_{2}\leq r^{2}(x_{3})\}$ and satisfying (3.13).
Proof.
It suffices to observe that $Dg$ depends continuously on $r^{\prime}$ and the estimates are independent of $r$ (they depend on the ratio $\frac{\rho}{r}$ which we keep constant). ∎
Lemma 3.4.
Let $R>0$ and let $T_{1},T_{2},\dots T_{k}$ be 3-simplexes such that $B_{\mathbb{R}^{3}}(0,2R)\subset\bigcup_{i=1}^{k}T_{i}=:G$ and $(0,0,0)$ is a common vertex of all $T_{i}$. Let $A_{i}$ be linear maps and
$f(x)=A_{i}x\quad x\in T_{i}$
be a sense-preserving piecewise affine homeomoprhism. Let the 2-subsimplexes of $T_{1},\dots,T_{k}$ containing $(0,0,0)$ be indexed as $F_{1},F_{2},\dots,F_{n}$ and the 1-subsimplexes containing $(0,0,0)$ be indexed as $S_{1},S_{2},\dots,S_{m}$. Then there exists an $r_{0}=r_{0}(R,\{S_{i}\}_{i=1}^{m})$ such that the sets $S_{i}+B(0,r_{0})\setminus B(0,R/2)$ are pairwise disjoint (and the dependence of $r_{0}$ on $R$ is linear). Let the numbers $r_{i}\leq r_{0}$ for $i=1,\dots,m$ and let the constants $w_{j}$ be so small that after an appropriate rotation in the preimage they satisfy (3.10) for all $x\in F_{j}\supset S_{i}$ such that $\operatorname{dist}(x,S_{i})>r_{i}/2$. Let $\hat{g}=\hat{g}_{\{r_{i}\_{i=1}^{m},\{w_{j}\}_{j=1}^{n}}$ be the diffeomorphism of $B(0,2R)\setminus B(0,R/2)$ constructed using Lemma 3.1 close to all $F_{j}\setminus S_{i}+B(0,r_{i}/2)$ and Lemma 3.2 close to all $S_{i}$. Then there exists an $\alpha,\beta>0$ such that if
$\frac{w_{j}}{R}\leq\alpha\quad\text{and}\quad\frac{r_{i}}{R}\leq\beta,$
and a diffeomorphism $g$ of $B(0,2R)$ such that $g=\hat{g}$ on $B(0,2R)\setminus B(0,R)$.
Proof.
Step 1. Projection to the sphere is injective. Let us prove that the map $x\to\frac{\hat{g}(x)}{|\hat{g}(x)|}$ is injective on $\partial B(0,t)$ for $t\in[R/2,2R]$. The claim is easy to observe when we replace $\hat{g}$ with $f$ since $f$ is a piecewise affine homeomorphism. This means (since we assume that $f$ is sense-preserving) that $\det Df\geq C>0$, a.e. on $G$. For each point $x$ in the preimage we consider an orthonormal positively-oriented basis of $\mathbb{R}^{3}$$\{\frac{x}{|x|},u_{2},u_{3}\}$. For each point $f(x)$ in the image we consider an orthonormal positively-oriented basis of $\mathbb{R}^{3}$$\{\frac{f(x)}{|f(x)|},v_{2},v_{3}\}$. We calculate $J_{f}$ with respect to this basis. Since $\partial_{\frac{x}{|x|}}f=\frac{f(x)}{|x|}$ we get that the subdeterminant of $Df$ with respect to $\{u_{2},u_{3}\}\times\{v_{2},v_{3}\}$ is necessarily positive and by linearity is independent of $|x|$. This implies that the map $x\to\frac{f(tx)}{|f(tx)|}$ defined on $\partial B(0,1)$ is a positively oriented bi-Lipschitz map of the unit sphere onto itself (and is independent of the parameter $t$).
We want to prove the same estimate for $\hat{g}$ as we have for $f$, i.e. the subdeterminant of $D\hat{g}$ with respect to $\{u_{2},u_{3}\}\times\{v_{2},v_{3}\}$ is positive for every $t$. We may assume that $J_{\hat{g}}\geq C>0$ on $B(0,2R)\setminus B(0,R/2)$. On the most of $\partial B(0,t)$ we are far away from any $F_{j}$ and so $\hat{g}=f$ (see (3.2)) and therefore we have the estimate immediately. Let us assume that we are close to some $F_{j}$ but not close to any $S_{i}$. By basic geometry and (3.3), we have that
$\Big{|}\partial_{\frac{x}{|x|}}\hat{g}-\partial_{\frac{\pi_{F_{j}}(x)}{|(x)|}}\hat{g}\Big{|}=\Big{|}\partial_{\frac{x-\pi_{F_{j}}(x)}{|x|}}\hat{g}\Big{|}\leq\frac{w_{j}}{R}|\partial_{v}\hat{g}|\leq C\frac{w_{j}}{R}\|Df\|_{L^{\infty}},$
where $\pi_{F_{j}}$ is the orthogonal projection of $\mathbb{R}^{3}$ onto $F_{j}$ and $v$ is a unit vector perpendicular to $F_{j}$. We calculate $J_{\hat{g}}$ by an expansion of the row corresponding to the component parallel to $\hat{g}(x)$ as
$J_{\hat{g}}\leq|D_{\frac{x}{|x|}}\hat{g}|\Delta+\big{|}D_{\frac{x}{|x|}}\hat{g}-D_{\frac{\pi_{F_{j}}(x)}{|(x)|}}\hat{g}\big{|}\|Df\|_{L^{\infty}}^{2}$
where by $\Delta$ we denote the subdeterminant of $D\hat{g}$ with respect to $\{u_{2},u_{3}\}\times\{v_{2},v_{3}\}$. This means that
$\Delta\geq|D_{\frac{x}{|x|}}\hat{g}|^{-1}\Big{(}J_{\hat{g}}-C\frac{w_{j}}{R}\|Df\|_{L^{\infty}}^{3}\Big{)}.$
Assuming that the ratio $\frac{w_{j}}{R}$ is sufficiently small we have that $\Delta$ is positive bounded away from zero.
Using the same argument on neighborhoods of $S_{i}$ we observe that also here $\Delta$ is positive bounded away from zero as soon as $\frac{r_{i}}{R}$ is sufficiently small. This implies that $\frac{\hat{g}}{|\hat{g}|}$ is a local homeomorphism of $\partial B(0,t)$ onto $\partial B(0,1)$, for $t\in[R/2,2R]$ and therefore, by Theorem 2.5, $\hat{g}$ is a homeomorphism.
Step 2. An injective convex combination of $\hat{g}$ and its projection onto a sphere. It follows from step 3 that the image of $B(0,\frac{3}{4}R)$ under $\hat{g}$ is a star-shaped domain containing the origin and having smooth boundary. This means we find a $\rho=C(f)r>0$ such that $B(0,2\rho)\subset\hat{g}(B(0,\frac{3}{4}R))$. We define $g$ on $B(0,R)\setminus B(0,\frac{3}{4}R)$ as follows
$g(x)=\eta\Big{(}\frac{4|x|-3r}{r}\Big{)}\hat{g}(x)+\Big{(}1-\eta\Big{(}\frac{4|x|-3r}{r}\Big{)}\Big{)}\rho|x|\frac{\hat{g}(x)}{|\hat{g}(x)|}.$
The injectivity of this map follows from the injectivity proved in the previous step and the choice of $\rho$, which guarantees that
$\Big{\langle}D_{\frac{x}{|x|}}\hat{g},\frac{\hat{g}(x)}{|\hat{g}(x)|}\Big{\rangle}>0.$
It follows that $J_{g}$ is bounded from below away from zero.
Step 3. Reduce to a linear map on $B(0,\tfrac{3}{4}R)\setminus B(0,\tfrac{1}{2}R)$. We have that the map $\frac{4}{3\rho}g(\frac{4}{3R}x)$ is a sense-preserving diffeomorphism of $S^{2}$ onto $S^{2}$. Then we may apply Theorem 2.2, to get a isotopy $\Psi:S^{2}\times[0,1]\to S^{2}$ such that $\Psi(\cdot,t)$ is a diffeomorphism of $S^{2}$ onto $S^{2}$, $\Psi(x,1)=\frac{4}{3\rho}g(\frac{4}{3R}x)$ and $\Psi(x,1)=x$. Then we may define
$g(x):=\rho|x|\Psi\Big{(}x,\eta\Big{(}\frac{4|x|-2R}{R}\Big{)}\Big{)}$
for $x\in B(0,\tfrac{3}{4}R)$. One may easily observe that $g(x)=\rho x$ on $B(0,R/2)$. It is easy to observe that $g$ is now a diffeomorphic extension of $\hat{g}$ restricted to $B(0,2R)\setminus B(0,R)$. ∎
3.1. Proof of Theorem A
Without loss of generality we may assume that $f$ is sense preserving (if not consider a reflection).
Let us first conduct the proof in a simplified case, i.e. where $f$ is a finitely piecewise affine homeomorhpism. We are able to resolve this case by choosing a constant $w_{j}$ for each 2-subsimplex $F_{j}$ and a constant $r_{i}$ for each 1-subsimplex $S_{i}$. Let the vertexes of the simplicial complex, i.e. its 0-sunsimplexes be indexed by $V_{k}$. We start by choosing a preliminary $\tilde{R_{k}}$ for each $V_{k}$ satisfying only the requirement that $B(V_{k},\tilde{R}_{k})$ are pairwise disjoint and that if $T_{m}\cap B(V_{k},\tilde{R}_{k})\neq\emptyset$, then $V_{k}\in T_{m}$ for each simplex $T_{m}$. Now we may choose $\tilde{r}_{i},\tilde{w}_{j}$ so small that the hypothesis of Lemma 3.2 and Lemma 3.4 are satisfied. Now we use Lemma 3.1, Lemma 3.2 and Lemma 3.4 to construct a diffeomorphism $\tilde{g}$ equaling $f$ whenever we are outside the set
$E_{1}:=\bigcup_{k}B(V_{k},\tilde{R}_{k})\cup\bigcup_{i}(S_{i}+B(0,\tilde{r}_{i}))\cup\bigcup_{j}(F_{j}+B(0,\tilde{w}_{j})).$
We need to refine this map, since we have no estimate on $|D\tilde{g}|$ on $B(V_{k},\tilde{R}_{k})$. Let us denote $M:=\|Dg\|_{L^{\infty}}$. We easily observe that $M$ is a function of the ratios $\frac{r_{i}}{R}$, $\frac{w_{j}}{R}$ and $\frac{w_{j}}{r_{i}}$, i.e. if we scale all of these simultaneously by some $\lambda\in(0,1]$ we get a diffeomorphism $g_{\lambda}$ each time also satisfying the estimate
(3.17) $\|Dg_{\lambda}\|_{L^{\infty}}=M.$
Let us denote by $K:=\|Dg^{-1}\|_{L^{\infty}}$, arguing in the same way we have:
(3.18) $\|Dg^{-1}_{\lambda}\|_{L^{\infty}}=K.$
On the other hand the measure of the set $E_{\lambda}=\{g_{\lambda}\neq f\}$ tends to zero as $\lambda\to 0^{+}$. By this argument we estimate that
$\|Df-Dg_{\lambda}\|_{L^{p}}^{p}\leq M^{p}\mathcal{L}^{3}(E_{\lambda}),$
for each $p\in[1,\infty)$.
We may estimate the error for the inverse in the same way, i.e. by the uniform Lipschitz quality of all $g_{\lambda}$ we have that
$\mathcal{L}^{3}(g_{\lambda}(E_{\lambda}))\leq M^{3}\mathcal{L}^{3}(E_{\lambda}).$
Further because $|Dg^{-1}_{\lambda}|$ is bounded independently of $\lambda$ we have that
$\|Df^{-1}-Dg^{-1}_{\lambda}\|_{L^{p}}\to 0\text{ as }\lambda\to 0$
for each $p\in[1,\infty)$.
Now we consider the general case where we have a locally finite simplicial complex. Let us choose an increasing sets sequence $A_{1}\subset A_{2}\subset\dots$, such that each $A_{n}$ is the finite union of simplexes $\{T_{k}\}_{k=1}^{N_{n}}$ and such that if $T_{k}\cap\partial A_{n}\neq\emptyset$ then $T_{k}\cap\partial A_{n+1}=\emptyset$. By $n$-border simplexes we mean those simplexes in $A_{n+1}\setminus A_{n}$ such that $T_{k}\cap\partial A_{n}\neq\emptyset$. By $H_{n}$ denote the set of simplexes in $A_{n}\setminus A_{n-1}$ and the $n$-border simplexes. For each $n$ we choose a value $\tilde{\lambda}_{n}$ so small that whenever we construct $\tilde{g}_{n}$ with $0<\lambda_{n}\leq\tilde{\lambda}_{n}$using the above process, we have
$\|f-\tilde{g}_{n}\|_{W^{1,p}}+\|f^{-1}-\tilde{g}^{-1}_{n}\|_{W^{1,q}}<2^{-n}\varepsilon,$
where the norms are taken over the interior of the union of simplexes of $H_{n}$ and the intersection of their images in $f$ and $\tilde{g}_{n}$ respectively.
In order to connect up the various $\tilde{g}_{n}$ into one diffeomorphism there is one more step that we need to conduct, namely on $n$-border simplexes we have defined $\tilde{g}_{n}$ and $\tilde{g}_{n+1}$. The idea is to allow $w_{j}$ and $r_{i}$ to smoothly change between the two values on the $F_{j}$ and $S_{i}$ respectively. On the other hand, if $F_{j}$ or $S_{i}$ are too short this may not be possible since in order for Lemma 3.1, Lemma 3.2 and Corollary 3.3 to hold we need $|Dw_{j}|$ and $|r^{\prime}|$ to be very small. On the other hand, now knowing the length of the segments $S_{i}$ we may now choose $\lambda_{n}<\tilde{\lambda}_{n}$ and $\lambda_{n+1}<\tilde{\lambda}_{n+1}$ so that the difference in respective values of $w_{j}^{n},w_{j}^{n+1}$ and of $r_{i}^{n},r_{i}^{n+1}$ are so small that there exist smooth functions $w_{j}$ and $r_{i}$ having the chosen values for $H_{n}$ near $A_{n}$ and having the values chosen for $H_{n+1}$ further away from $A_{n}$, while $|Dw_{i}|$ and $|r^{\prime}|$ being as small as required.
By the preceding lemmas (and Corollary 3.3) we construct a diffeomorphism $g$. By the choices of $\lambda_{n}<\tilde{\lambda}_{n}$ we have that
$\|f-g\|_{W^{1,p}}+\|f^{-1}-g^{-1}\|_{W^{1,q}}<\varepsilon,$
where the norms are taken over $\Omega$ and $f(\Omega)$ respectively.
3.2. Proof of Theorem B
By (3.17) and (3.18), for each $f$, there exists a $C(f)$ such that
$\|Df-Dg_{n}\|_{L^{\infty}(H_{n})}+\|Df^{-1}-Dg_{n}^{-1}\|_{L^{\infty}(f(H_{n}))}\leq C(f).$
Then using the fact that $\mathcal{L}^{3}(E_{\lambda_{n},n})\to 0$ as $\lambda_{n}\to 0$, where $E_{\lambda_{n}}=H_{n}\cap\{g_{\lambda_{n},n}\neq f\}$, $f$ is bi-Lipschitz on $H_{n}$ and Lemma 2.7 we see that
$\|Df-Dg_{n}\|_{X(H_{n})}+\|Df^{-1}-Dg_{n}^{-1}\|_{Y(f(H_{n}))}\to 0$
as $\lambda_{n}\to 0$. The estimates for the function values are even easier.
4. Remark on the proof for higher dimensions
The proof in dimensions $d=4$ follows almost exactly the same way as the case in $d=3$, since the arguments in Lemma 3.1 are independent on the number of dimensions of the hyperplane $\{x_{1}=0\}$ and the proof of Lemma 3.2 does not change when we replace the segment in $\mathbb{R}^{3}$ with a 2 dimensional plane in $\mathbb{R}^{4}$. The only notable difference is that instead of Theorem 2.2 we use the following generalization:
Theorem 4.1.
For $d=2,3,4,6,12,56,61$ and every $\alpha_{0}$, $\mathcal{C}^{\infty}$, a sense-preserving diffeomorphism of $S^{d-1}=\partial B_{\mathbb{R}^{d}}(0,1)$ onto itself there exists an $\alpha\in\mathcal{C}^{\infty}([0,1]\times S^{d-1},S^{d-1})$ such that $\alpha(t,\cdot)$ is a diffeomorphism of $S^{d-1}$ for each $t$, $\alpha(0,\cdot)=\alpha_{0}(\cdot)$ and $\alpha(1,\cdot)=\operatorname{id}(\cdot)$.
This theorem is equivalent to stating that the group of sense-preserving diffeomorphisms of the sphere $S^{d-1}$ is connected for $d=2,3,4,6,12,56,61.$ As mentioned above for $d=3$ the theorem was proven by Smale in [24, Theorem A]. For $d=4$ it was proven by Hatcher in [10] and the cases of $d\geq 6$ can be found in [9, Section II] and [13, Table 1 and 2], where can also be found that the theorem fails for all other $d\leq 90$ outside $d=5.$ It is still an open problem for $d=5$ and whether it holds for any $d>90.$
Since the method of our proof is contingent on finding such an isotopy $\alpha$, the obstacle for generalizing to higher dimension than $d=4$, is that we do not know that the space of sense-preserving diffeomorphisms from the sphere onto itself is connected for dimension $5$ and in fact is false for dimensions other than the ones mentioned in Theorem 4.1.